# PACP: Priority-Aware Collaborative Perception for Connected and Autonomous Vehicles

### v2

@@ 2.1 Background of Two Schemes
(1) Fairness-based perception scheme: For perception data from different sources, achieving fairness is a popular design goal[36], which ensures that equal resources for different resources to transmit perception data in V2V networks. The Jain’s fairness index $J$ is the most popular metric to characterize the fairness[33]:
@@
where $n$ is the total number of nodes and $x_{i}$ is the resource allocated to the $i$-th node. According to Cauchy-Schwartz inequality, its value ranges from $(0,1]$, which is equal to 1 if and only if all $x_{i}$ are equal, i.e., 1 indicates perfect fairness. Based on the type of resources, two popular fairness schemes are the subchannel- and throughput-fairness schemes, respectively. Specifically, the subchannel-fairness scheme means each CAV can be allocated the same amount of spectrum resources while the throughput-fairness scheme requires each CAV to have the same transmission rate.
(2) Priority-aware perception scheme: Traditional fairness-based schemes treat all CAVs equally, potentially enhancing the quality of less important information while diminishing the quality of crucial data, consequently reducing the efficiency of the system. Therefore, it is crucial to determine the priority of the sensed data from the nearby CAVs. Existing works have investigated several popular priority factors as priority, such as link latency[37] and routing situation[38]. However, previous studies failed to investigate the negative effects of a “poisonous” CAV, which transmits its perception result to ego CAV but aggravates the performance of collaborative perception because of poor channel condition and specific location. Therefore, the ego vehicle gives higher priority to the vehicles with more important data and better BEV match (refer to details in Sec. 4.2). In so doing, the ego CAV assigns a lower priority to those poisonous CAVs, mitigating the adverse impact on AP@IoU.
2.2 Comparison Between Fairness and Priority
@@
To better understand the limitations of fairness-based schemes, we use two typical fairness-based schemes to showcase their resource allocation and BEV prediction in comparison to the priority-aware scheme. Fig. 1 shows the bandwidth and throughput allocation of three schemes in a V2V network. In our scenario, the ego CAV exploits three schemes to combine three nearby CAVs’ camera perception results for BEV prediction, respectively. The whole bandwidth is 200 MHz.
In the subchannel-fairness scheme (Fig. 1(a)), each CAV is allocated an equal amount of spectrum resources, which can lead to inefficient resource utilization, especially when some CAVs have poor channel conditions. The throughput-fairness scheme (Fig. 1(b)) aims to equalize the transmission rates among CAVs by compensating for weaker channels with more resources. However, this approach can still result in suboptimal BEV prediction accuracy. Specifically, if a CAV’s channel quality is extremely poor (e.g., CAV 1), the allocated discrete bandwidth might still be insufficient to ensure equalized throughput, leading to reduced perception performance.
In contrast, the priority-aware perception scheme (Fig. 1(c)) dynamically adjusts the priority weights of each CAV based on channel resources, perception accuracy, and perception coverage. Due to the worst channel condition, it is noted that CAV 1 has the lowest priority, thereby its data is discarded by the priority-aware scheme. This approach ensures that the most critical and high-quality data from the CAVs with the best channel conditions and the closest proximity to the ego vehicle are transmitted with minimum loss and latency. This selective collaboration leads to a significant improvement in BEV predictions. Fig. 2(b) shows the impact of different communication conditions on the dynamic AP@IoU (mobile object detection). The superiority of the priority-aware scheme becomes evident, as the dynamic AP@IoU rises to 0.685. Therefore, the priority-aware scheme surpasses the fairness-based methods by discerningly choosing which CAVs to collaborate with, leading to improved BEV predictions.
@@ 3 System Model
$K$ The number of orthogonal sub-channels
$N$ The total number of nodes or vehicles in the network
@@
$\mathcal{D}$ The matrix of transmission rates, $\mathcal{D}=\{d_{ij}\}_{N\times N}$
$r_{ij}$ The adaptive compression ratio for data transmission between CAVs $i$ and $j$
$E_{ij}^{t}$ The energy consumption bounds for data transmission between CAVs $i$ and $j$
$E_{j}^{c}$ The energy consumption bounds for computation at CAV $j$
$\mathcal{P}_{ij}$ The priority weight between the ego vehicle $j$ and nearby CAV $i$
@@
$\omega_{1},\omega_{2}$ Weights for perception quality and perceptual region, respectively
In this section, we present a V2X-aided collaborative perception system with CAVs, including the system’s structure, channel modeling, the constraints of computational capacity and energy, and background information on IoU. The key notations are listed in the Table II.
3.1 System Overview
Figure 3: Overview of V2X-aided collaborative perception system.
3.2 Channel Modeling and System Constraints
@@
Let $\mathcal{D}=\left\{d_{ij}\right\}_{N\times N}$ be the matrix consisting of transmission rates, where each element is non-negative, for $\forall(i,j)\in\mathbf{E}$. Each element $d_{ij}$ represents the data rate from vehicle $v_{i}$ to vehicle $v_{j}$ and is subsequently processed by $v_{j}$. It is noteworthy that $d_{ij}$ satisfies:
${r}_{ij}d_{ij}\leq\min(C_{ij},A_{i}),$ (3)
Here, ${r}_{ij}\in(0,1]$ denotes the adaptive compression ratio, obtained by the compression algorithm outlined in Sec. 5.4. $A_{i}$ signifies the amount of local perception data at $v_{i}$ per second, i.e., perception data generation rate at the location of $v_{i}$. This constraint implies that the value of $d_{ij}$ must be limited either by the achievable data rate or by the rate of perception of the local data present at vehicle $v_{i}$. Furthermore, it is observed that an inadequate compression ratio results in a diminution in the accuracy of perception data, while an excessively high compression ratio results in suboptimal throughput maximization. Consequently, the constraint for the compression ratio is defined as follows:
@@
3.3 The Background Information on IoU
(a) The intersection of $G$ and $P$.
(b) The union of $G$ and $P$.
The respective areas of $G$ and $P$ are defined as follows:
$S_{G}$ $=|(x_{0}^{G}-x_{1}^{G})\times(y_{0}^{G}-y_{1}^{G})|,$ (10)
$S_{P}$ $=|(x_{0}^{P}-x_{1}^{P})\times(y_{0}^{P}-y_{1}^{P})|.$
$S_{G\cap P}=\left|(x_{0}^{I}-x_{1}^{I})\times(y_{0}^{I}-y_{1}^{I})\right|$ (11)
$S_{G\cup P}=S_{G}+S_{P}-S_{G\cap P}$ (12)
IoU is thus calculated as:
@@ 4.1 Selection of Priority Weights
There exist many ways to define priority weights, such as distance[40], channel state[41], and information redundancy[42]. However, as CoBEVT is the backbone for RGB data fusion[7], our priority weight definition is based on the IoU of BEV features between adjacent CAVs. CoBEVT’s local attention aids in pixel-to-pixel correspondence during object detection fusion. IoU of the BEV features reveals the consistency in environmental perception between CAVs. Factors like channel interference and network congestion may cause inconsistencies. As consistency is crucial in fusion data, inconsistencies can lead to data misrepresentations. Hence, we design a BEV-match mechanism by relying on IoU analysis in the next subsection, giving preference to CAVs with closely aligned perceptions to the ego CAV.
@@ 4.2 Procedure of Obtaining Priority Weights
Figure 6: The example of calculating priority weight $\mathcal{P}_{10}$.
@@ 5.1 Nonlinear Programming Problem Analysis
For $j=1,2,\ldots,N$, the sub-problem $\mathbf{P}_{1}$ is an NLP problem due to the nonlinear constraints given by (3), (21a), and (21b) [^3] This is due to the product of decision variables ${r}_{ij}$ and $d_{ij}$.. While global optimization techniques like branch and bound or genetic algorithms can be used for non-convex problems, they are computationally demanding. Our approach is to linearize the problem. We define $\mathbf{U}=\mathcal{R}\odot\mathcal{D}=\left[u_{ij}\right]_{N\times N}$ where $\odot$ is the Hadamard product and $u_{ij}={r}_{ij}d_{ij}$. This linearizes the product term in the constraints. Thus, $\mathbf{P}_{1}$ is reformulated as [^4] For simplicity, we omit $\mathcal{I}\left(\mathcal{S}_{i}\right)$, $\omega_{1}$, and $\omega_{2}$ in the formulation of the first sub-problem, since those terms are independent of $\mathbf{P_{1-1}}$.:
@@
$\frac{u_{ij}}{{r}_{j,\max}}\leq d_{ij}\leq\frac{u_{ij}}{\max\left({r}_{j,\min},\frac{\eta}{e^{L_{ij}}}\right)},$ (23)
@@
$\mathbf{P_{1-2}}$ is a standard linear programming solvable using techniques like the simplex or interior-point methods. If the optimal outcome of $\mathbf{P_{1-2}}$ is $u_{ij}^{(n)}$, the optimal solutions for transmission rate and adaptive compression ratio are $d_{ij}^{(n)}={u_{ij}^{(n)}}\left[{{\max\left({r}_{j,\min},\frac{\eta}{e^{L_{ij}}}\right)}}^{-1}\right]$ and ${r}_{ij}^{(n)}={u_{ij}^{(n)}}/{d_{ij}^{(n)}}$, respectively. Considering that $d_{ij}$ values are at the edges of the feasible region, the optimal solution for $\mathbf{P_{1-2}}$ matches that of $\mathbf{P_{1-1}}$. It is noted that $u_{ij}=0\ \text{if}\ s_{ij}^{(n-1)}=0$ guarantees that $u_{ij}$ is explicitly set to zero whenever $s_{ij}^{(n-1)}=0$.
@@ 5.3 Submodular Analysis and Solutions
Similarity,
$f(\mathcal{S}_{i})=g(\mathcal{S}_{i})+h(\mathcal{S}_{i}),$ (27)
@@ 5.4 Deep Learning-Based Adaptive Compression
where function $G$ is the R-D mapping function, which utilizes a lookup table and converts the compression ratio to the tradeoff parameter $\beta$. Therefore, there is no explicit expression for the function $G$. Moreover, the adaptive compression network is based on a pre-trained model[28], using historical camera data as training dataset to obtain network parameter $\theta,\xi$ by solving the multi-R-D problem in (34).
1) Encoder Phase: The input image $\mathbf{x}$ undergoes transformation to latent representation $\mathbf{z}$. A quantization function ${Q}$ further refines $\mathbf{z}$ into a discrete vector $q\in\mathbb{H}^{D}$, leading to $q=Q(\mathbf{z})$. In V2V scenarios, $q$ is serialized into bitstream $b$, with entropy coding to maximize the bandwidth efficiency. As shown in Fig. 8(a), the encoder compresses the input camera data, reducing its size for transmission while retaining essential information. It comprises the following layers:
(a) The encoder part.
(b) The decoder part.
2) Decoding Phase: The bitstream $b$ undergoes inverse transformations. This includes dequantization ($\hat{\mathbf{z}}=Q^{-1}(q)$) and a reconstruction function ($\hat{\mathbf{x}}=H({\hat{\mathbf{z}}})$) to retrieve the original image. As shown in Fig. 8(b), the decoder reconstructs the original camera data from the compressed bitstream, ensuring that the essential information is preserved and accurately recovered. It consists of:
Figure 9: The architecture of modulation/ demodulation networks.
3) Modulation/Demodulation Networks: As shown in Fig. 9, the modulation and demodulation networks translate the CSI feedback into adaptive modulation and demodulation vectors, respectively, for the encoder and decoder. They include:
4) Fine-tuning Phase: This adaptive approach enables the ego CAV and nearby CAVs to adaptively adjust the data compression ratio according to dynamic channel conditions. As shown in Fig. 7, we introduce a fine-tuning compression strategy to further reduce temporal redundancy between frames. Utilizing a Modulated Autoencoder (MAE), our method first transmits a few uncompressed images to RSU for real-time fine-tuning.
@@
0:  Priority weight $\mathcal{P}$, near-optimal compression ratio $\mathcal{R}$, link establishment $\mathcal{S}$, data rate $\mathcal{D}$, modulated autoencoder, and BEV prediction.
1:  Calculate priority weight $\mathcal{P}$ according to Sec. 4;
2:  Initialize the link establishment decision $\mathcal{S}^{(0)}$;
@@
8:     Solve the submodular problem $\mathbf{P_{2}}$ using Algorithm 1;
9:  end while
10:  Obtain the near-optimal solution for $\mathcal{R},\mathcal{S},\mathcal{D}$;
11:  Determine the trade-off parameter $\beta^{*}$ based on the channel state;
12:  Train encoders and decoders to compress and reconstruct raw camera data according to Eq. (34);
13:  Predict BEV feature using the reconstructed camera data.
@@ 6.3 Analysis of Experimental Results
Intuitively, we can adopt throughput as a metric for cooperative perception. Similar to utility value, a higher throughput often entails a greater amount of valuable information being transmitted. Fig. 12(a) shows that larger power enhances throughput. Fig. 12(b) depicts an inverse correlation between the vehicle distribution maximum range (80 m - 210 m) and throughput. It is noted that the performance of FTS strategy deteriorates because of balancing communication channels of remote vehicles. Remarkably, with five cooperative CAVs, the PACP algorithm augments the average throughput by 20.39% compared with DMDDA. Furthermore, Fig. 12(d) underscores the linear affinity between the network’s bandwidth (80-180 MHz) and throughput. Specifically, a bandwidth of $W=180$ MHz elevates the average throughput by 18.61% using the PACP algorithm against other baselines. To sum up, it is evident that a higher throughput results in a richer amount of high-priority data being transmitted according to Figs. 11-12.

### v3

@@ 2.1 Background of Two Schemes
Fairness-based scheme: This scheme aims to achieve fairness in resource allocation among different CAVs. The Jain’s fairness index is used to measure fairness:[33]:
@@
where $n$ is the total number of nodes and $x_{i}$ is the resource allocated to the $i$-th node. A perfect fairness index of 1 indicates equal resource allocation. Two common fairness schemes are subchannel-fairness (equal spectrum resources) and throughput-fairness (equal transmission rates).
Priority-aware scheme: Unlike fairness-based schemes, this scheme assigns different priority levels to CAVs based on the importance and quality of their data. The ego vehicle gives higher priority to CAVs with better channel conditions and more crucial perception data. Existing works have investigated several popular priority factors as priority, such as link latency[36] and routing situation[37]. This approach mitigates the negative impact of "poisonous" CAVs with poor channel conditions or less relevant data, thus enhancing the overall system efficiency. Compared to the fairness-based scheme, the advantages of priority-aware perception can be summarized as follows:
  • Transmission Efficiency: By prioritizing CAVs with better data quality and channel conditions, the priority-aware scheme optimally allocates spectrum resources, ensuring efficient transmission.
  • Improved Prediction: This scheme reduces the influence of poisonous data, leading to more accurate BEV predictions by focusing on high-quality perception inputs.
  • Dynamic Adaptability: The priority-aware scheme dynamically adjusts to changing channel conditions and task requirements, maintaining robust performance in diverse environments.
2.2 An Illustrative Motivating Example
@@
To understand the limitations of fairness-based schemes, we compare their resource allocation and BEV prediction against an ad hoc priority-aware scheme. Fig. 1 illustrates the bandwidth and throughput allocation of three schemes in a V2V network, with a total bandwidth of 200 MHz.
In the subchannel-fairness scheme (Fig. 1(a)), each CAV receives an equal amount of spectrum resources, leading to inefficient utilization, especially with poor channel conditions. The throughput-fairness scheme (Fig. 1(b)) equalizes transmission rates by allocating more resources to weaker channels, but it can still result in suboptimal BEV predictions. For instance, if a CAV has extremely poor channel quality (e.g., CAV 1), even the allocated bandwidth may be insufficient, reducing perception performance.
In contrast, the priority-aware perception scheme (Fig. 1(c)) dynamically adjusts the priority weights of each CAV based on channel resources, perception accuracy, and coverage. For example, CAV 1, with the worst channel condition, is assigned with the lowest priority and its data is discarded. This approach ensures that critical and high-quality data from CAVs with the best channel conditions and/or proximity to the ego vehicle are transmitted with minimal loss and latency, significantly improving BEV predictions. Fig. 2(b) demonstrates the superior performance of the priority-aware scheme, with the dynamic AP@IoU metric rising to 0.685. Thus, this ad hoc priority-aware scheme outperforms fairness-based methods by selectively collaborating with the most valuable CAVs, leading to enhanced BEV predictions. However, how to design an effective priority-aware scheme is critical, which motivates this research.
@@ 3 System Model
In this section, we present a V2X-aided collaborative perception system with CAVs, including the system’s structure, channel modeling, and the constraints of computational capacity and energy. The key notations are listed in the Table II.
3.1 System Overview
Figure 3: Overview of V2X-aided collaborative perception system.
3.2 Channel Modeling and System Constraints
@@
$N$ The total number of vehicles in the network
$K$ The number of orthogonal sub-channels for the whole V2V networks
$W$ The total bandwidth of the V2V networks
@@
$\mathcal{D}$ The matrix of transmission rates
$r_{ij}$ The adaptive compression ratio
$E_{ij}^{t}$ The energy bounds for data transmission
$E_{j}^{c}$ The energy consumption bounds for computation
$E_{j}^{T}$ The energy consumption threshold
$P_{t}$ Transmission power of each CAV
$F_{j}$ CPU capacity of CAV $j$
$\beta$ Model complexity parameter depending on the architecture of the neural networks
$A_{j}$ Local data generation rate per second at CAV $j$
$\mathcal{P}_{ij}$ The priority weight between two CAVs
@@
$(\omega_{1},\omega_{2})$ Weights for perception quality and region
$(r_{\min},r_{\max})$ Compression ratio range
@@
Let $\mathcal{D}=\left\{d_{ij}\right\}_{N\times N}$ be the matrix consisting of transmission rates, where each element is non-negative, for $\forall(i,j)\in\mathbf{E}$. Each element $d_{ij}$ represents the data rate without compression from vehicle $v_{i}$ to vehicle $v_{j}$, which is then processed by $v_{j}$. It is noteworthy that $d_{ij}$ satisfies:
${r}_{ij}d_{ij}\leq\min(C_{ij},{r}_{ij}A_{i}),$ (3)
Here, ${r}_{ij}\in(0,1]$ denotes the adaptive compression ratio, obtained by the compression algorithm outlined in Sec. 5.4. $r_{ij}d_{ij}$ represents the actual transmission rate after compression. $A_{i}$ signifies the amount of local perception data at $v_{i}$ per second, i.e., perception data generation rate at the location of $v_{i}$. This constraint implies that the actual transmission rate must be limited either by the achievable data rate or by the locally compressed data present at vehicle $v_{i}$. Furthermore, it is observed that an inadequate compression ratio results in a diminution in the accuracy of perception data, while an excessively high compression ratio results in suboptimal throughput maximization. Consequently, the constraint for the compression ratio is defined as follows:
@@ 4.1 Selection of Priority Weights
There exist many ways to define priority weights, such as distance[39], channel state[40], and information redundancy[41]. However, as CoBEVT is the backbone for RGB data fusion[7], our priority weight definition is based on the Intersection over Union (IoU) of BEV features between adjacent CAVs. IoU is an effective metric in computer vision, especially in object detection, used to quantify the overlap between two areas. The background of IoU can be obtained in CoBEVT’s local attention aids in pixel-to-pixel correspondence during object detection fusion. IoU of the BEV features reveals the consistency in environmental perception between CAVs. Factors like channel interference and network congestion may cause inconsistencies. As consistency is crucial in fusion data, inconsistencies can lead to data misrepresentations. Hence, we design a BEV-match mechanism by relying on IoU analysis in the next subsection, giving preference to CAVs with closely aligned perceptions to the ego CAV.
@@ 4.2 Procedure of Obtaining Priority Weights
Figure 5: An example of determining priority weight $\mathcal{P}_{10}$.
@@ 5.1 Nonlinear Programming Problem Analysis
For $j=1,2,\ldots,N$, the sub-problem $\mathbf{P}_{1}$ is an NLP problem due to the nonlinear constraints given by (3), (15a), and (15b) [^3] This is due to the product of decision variables ${r}_{ij}$ and $d_{ij}$.. While global optimization techniques like branch and bound or genetic algorithms can be used for non-convex problems, they are computationally demanding. Our approach is to linearize the problem. We define $\mathbf{U}=\mathcal{R}\odot\mathcal{D}=\left[u_{ij}\right]_{N\times N}$ where $\odot$ is the Hadamard product and $u_{ij}={r}_{ij}d_{ij}$. This linearizes the product term in the constraints. Thus, $\mathbf{P}_{1}$ is equivalently reformulated as [^4] For simplicity, we omit $\mathcal{I}\left(\mathcal{S}_{i}\right)$, $\omega_{1}$, and $\omega_{2}$ in the formulation of the first sub-problem, since those terms are independent of $\mathbf{P_{1-1}}$.:
@@
$\frac{u_{ij}}{r_{j,\max}}\leq d_{ij}\leq\min\left\{A_{i},u_{ij}\left[\max\left(r_{j,\min},\frac{\eta}{e^{L_{ij}}}\right)\right]^{-1}\right\},$ (17)
@@
$\mathbf{P_{1-2}}$ is a standard linear programming solvable using techniques like the simplex or interior-point methods. If the optimal outcome of $\mathbf{P_{1-2}}$ is $u_{ij}^{(n)}$, the optimal solutions for transmission rate and adaptive compression ratio are $d_{ij}^{(n)}=\min\left\{A_{i},u_{ij}\left[\max\left(r_{j,\min},\frac{\eta}{e^{L_{ij}}}\right)\right]^{-1}\right\}$ and ${r}_{ij}^{(n)}={u_{ij}^{(n)}}/{d_{ij}^{(n)}}$, respectively. Considering that $d_{ij}$ values are at the edges of the feasible region, the optimal solution for $\mathbf{P_{1-2}}$ matches that of $\mathbf{P_{1-1}}$. It is noted that $u_{ij}=0\ \text{if}\ s_{ij}^{(n-1)}=0$ guarantees that $u_{ij}$ is explicitly set to zero whenever $s_{ij}^{(n-1)}=0$.
@@ 5.3 Submodular Analysis and Solutions
Please refer to Appendix C. $\blacksquare$
@@ 5.4 Deep Learning-Based Adaptive Compression
where function $G$ is the R-D mapping function, which utilizes a lookup table and converts the compression ratio to the tradeoff parameter $\beta$. Therefore, there is no explicit expression for the function $G$. Moreover, the adaptive compression network is based on a pre-trained model[28], using historical camera data as training dataset to obtain network parameter $\theta,\xi$ by solving the multi-R-D problem in (23). The detailed architecture and processes are described in Appendix D. We also evaluate the R-D performance of our proposed PACP framework using the Multi-scale Structural Similarity (MS-SSIM) and Peak Signal-to-Noise Ratio (PSNR) metrics in Appendix D.5.
@@
10:  Priority weight $\mathcal{P}$, near-optimal compression ratio $\mathcal{R}$, link establishment $\mathcal{S}$, data rate $\mathcal{D}$, modulated autoencoder, and BEV prediction.
1:  Initialize priority weight $\mathcal{P}^{(0)}$ by equally allocating bandwidth and transmitting initial frames;
22:  Initialize the link establishment decision $\mathcal{S}^{(0)}$;
@@
48:     Solve the submodular problem $\mathbf{P_{2}}$ using Algorithm 1;
59:     Update priority weight $\mathcal{P}$ based on the current bandwidth allocation and perception results;
10:     if Priority weight $\mathcal{P}$ changes exceed the threshold $\epsilon_{\mathcal{P}}$ then
11:        Re-initialize the link establishment decision $\mathcal{S}^{(0)}$, then re-calculate initial priority weight $\mathcal{P}$;
12:        Continue the optimization iteration for problem $\mathbf{P}$;
613:     end if
714:  end while
15:  Obtain the near-optimal solution for $\mathcal{R},\mathcal{S},\mathcal{D}$;
16:  Determine the trade-off parameter $\beta^{*}$ based on the channel state;
17:  Train encoders and decoders to compress and reconstruct raw camera data according to Eq. (23);
18:  Predict BEV feature using the reconstructed camera data.
@@ 6.3 Analysis of Experimental Results
Intuitively, we can adopt throughput as a metric for cooperative perception. Similar to utility value, a higher throughput often entails a greater amount of valuable information being transmitted. Fig. 9(a) shows that larger power enhances throughput. Fig. 9(b) depicts an inverse correlation between the vehicle distribution maximum range (80 m - 210 m) and throughput. It is noted that the performance of FTS strategy deteriorates because of balancing communication channels of remote vehicles. Remarkably, with five cooperative CAVs, the PACP algorithm augments the average throughput by 20.39% compared with DMDDA. Furthermore, Fig. 9(d) underscores the linear affinity between the network’s bandwidth (80-180 MHz) and throughput. Specifically, a bandwidth of $W=180$ MHz elevates the average throughput by 18.61% using the PACP algorithm against other baselines. To sum up, it is evident that a higher throughput results in a richer amount of high-priority data being transmitted according to Figs. 8-9.